A seller increased the quantity he offered for sale from 200 units to 250 units when price of his product increased by 12.5%. what is the price elasticity of supply of his product?


1
Expert's answer
2021-03-02T17:45:11-0500

The price elasticity of supply can be calculated as follows:


"E_s=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{Q_1}\times100\%}{\%\Delta P},""E_s=\dfrac{\dfrac{250-200}{200}\times100\%}{12.5\%}=2.0."